suppose 2.4 million lbs of almonds are sold when the price is $5.50 per pound and 4.8 million lbs of almonds are sold at $4.50 per pound. find a linear function, A(p), that expresses the amount of almonds sold as a function of the price per pound.
Answer by stanbon(75887)

suppose 2.4 million lbs of almonds are sold when the price is $5.50 per pound and 4.8 million lbs of almonds are sold at $4.50 per pound. find a linear function, A(p), that expresses the amount of almonds sold as a function of the price per pound.
--------
You have two points: (5.5,2.4 million) and (4.5,4.8 million)
----------------
slope = (4.8-2.4)/(4.5-5.5) = -2.4
-------------
equation form y = mx+ b
4.8 = -2.4*4.5 + b
b = 4.8 + 10.8
b = 15.6
-------------
A(p) = -2.4p +15.6 in millions of pounds of almonds.
=============
Cheers,
Stan H.